on the final exam in a large statistics class the average score was 76 with a standard deviation of 7. assuming a normal population, what scores were so deviant that their probability of occurrence was .05 or less? (hint: .025 in each tail)
Answer by stanbon(75887)

on the final exam in a large statistics class the average score was 76 with a standard deviation of 7. assuming a normal population, what scores were so deviant that their probability of occurrence was .05 or less? (hint: .025 in each tail)
-----
Find the z-value with a left tail of 2.5%
z = 1.96
Use x = z*s + u
-----
x = -1.96*7+76 = 62.28 (lower limit score)
x = +1.96*7+76 = 89.72 (upper limit score)
===========
Cheers,
Stan H.
===========